OpenStudy (anonymous):

Can somebody please help A system of linear inequalities is shown below: y ≥ -2x - 2 y + x ≥ 0 Describe the steps to graph the solution set to the system of inequalities. (10 points)

OpenStudy (cggurumanjunath):

Y>=-2X-2 Y+X>=0; Y<=-X -X>=-2X-2 -X+2X>=-2 X>=-2